Following are the American College Test (ACT) scores attained by the 25 members of a local high school graduating class.
21 24 23 17 31 19 19 20 19 25 17 23 16
21 20 28 25 25 21 14 19 17 18 28 20

Find P5, P10, and P20 for the ACT scores.

Answer by stanbon(75887)

Following are the American College Test (ACT) scores attained by the 25 members of a local high school graduating class.
21 24 23 17 31 19 19 20 19 25 17 23 16
21 20 28 25 25 21 14 19 17 18 28 20
Find P5, P10, and P20 for the ACT scores.
---------------------
I'll find P5.
---
1st: Put the data in ascending order from minimum to maximum
2nd: Find the locator for P5:
5%(25) = (5/100)25 = 5/4 = 1.25
Round up to the next whole number:
locator = 2
---
3rd Find P(5) which is the 2nd data element starting at the min.
----
P(5) = 16
===================
Cheers,
Stan H.